OpenStudy (anonymous):

Using the point-slope equation, find the equation containing (-5, -8) and slope m = -3

OpenStudy (anonymous):

Point slope form is y=y1+m(x-x1) where (x1,y1) is a point and m is your slope.. Plugging the values in: y=-8 -3(x+5)

OpenStudy (anonymous):

y+8=-3(x+5) {using formula y-y1=m(x-x1)} 3x+y +23=0

OpenStudy (anonymous):

You can further simplify by distributing the -3 through, and adding like terms. y=-8-3x-15 Add -8 and -15 together, -23 so y=-3x -23